OpenStudy (anonymous):

evaluate, (3/2)^4 = type an improper fraction? anyone

OpenStudy (jhannybean):

\[\large \left(\frac{3}{2}\right)^4\] First you want to distribute the 4 into the numerator and denominator. What do you get?

OpenStudy (anonymous):

12 over 8

OpenStudy (jhannybean):

\[\large \left(\frac{3^4}{2^4}\right)= \frac{(3)(3)(3)(3)}{(2)(2)(2)(2)}\]

OpenStudy (anonymous):

81/16 so im right.

OpenStudy (anonymous):

(3/2)^4 how do you convert this into mixed number??

OpenStudy (jhannybean):

Oh not just a fraction.... ah...hmm

OpenStudy (anonymous):

but im right on the first one the 81/16?

OpenStudy (jhannybean):

Yeah the 12/8 wasn't right though.

OpenStudy (anonymous):

how to convert that to mixed number

OpenStudy (jhannybean):

ok what if we had 80/16 what would that be?

OpenStudy (anonymous):

i think its 5 1/16, for the mixed number, im not sure if thats right?

OpenStudy (jhannybean):

That is correct :)

OpenStudy (jhannybean):

http://www.sketchtoy.com/41296567

OpenStudy (anonymous):

okay thanks

OpenStudy (jhannybean):

np